# Optimal frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v1

@@ 3.1. The Relationship between Tight Frames and Convex Polygons in $\mathbb{R}^{2}$
(ii) If $\bm{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfies (1.11), by Theorem 1.3, there exists an optimal polygon $P\in{\mathcal{E}}_{m}$ with the edge vectors $\{\varepsilon_{j}(\cos(\frac{j}{m}\pi),\sin(\frac{j}{m}\pi))^{T}:j\in[m]\}$. Note that for $j\in[m]$ we have
$f({\bm{\nu}}_{j})=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}\quad\text{and}\quad f({\bm{\nu}}_{j}^{\perp})=-(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}.$ (4.13)
@@
Hence, the edge set of $P$ can be rewritten as $f({\bm{\nu}}_{j}):j\in I\}\cup\{f({\bm{\nu}}_{j}^{\perp}):j\in[m]\backslash I\}$ where $I=\{j\in[m]:\varepsilon_{j}=1\}$, and such a polygon $P$ is unique up to translation. Then it follows from Theorem 1.2 that the set ${\bm{\nu}}_{j}:j\in I\}\cup\{{\bm{\nu}}_{j}^{\perp}:j\in[m]\backslash I\}$ is an optimal tight frame.
@@
(iii) Let ${\mathcal{A}}=\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\in\mathcal{T}_{m}$. As established in the proof of (i), $\mathcal{A}$ is an optimal tight frame. Hence, by Theorem 1.2, there is an optimal strictly convex $m$-gon $P$ with the edge set $\{f(\bm{a}_{1}),\dotsc,f(\bm{a}_{m})\}$. According to Theorem 1.3, there exists a vector $\bm{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfying (1.11), such that after a proper scaling and rotation, the edge set $\{f(\bm{a}_{1}),\dotsc,f(\bm{a}_{m})\}$ equals $(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}:j\in I\}\cup\{-(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}:j\in[m]\backslash I\}$, where $I=\{j\in[m]:\varepsilon_{j}=1\}$. By (4.13), this edge set can be rewritten as $f({\bm{\nu}}_{j}):j\in I\}\cup\{f({\bm{\nu}}_{j}^{\perp}):j\in[m]\backslash I\}$. Note that the function $f(\cdot)$ is invertible on $\mathcal{T}$. It follows that ${\mathcal{A}}=\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}=\{{\bm{\nu}}_{j}:j\in I\}\cup\{{\bm{\nu}}_{j}^{\perp}:j\in[m]\backslash I\}$, up to scaling and rotation. This completes the proof. ∎
@@ 5. Discussion and Future Work
We conclude this paper by highlighting an intriguing connection between the perimeter-maximizing polygon problem and a fundamental question in discrepancy theory. Assume that $E=\{\bm{e}_{1},\dotsc,\bm{e}_{m}\}$ is the edge set of a convex $m$-gon ${P}\subset\mathbb{R}^{2}$ with fixed perimeter $\mathrm{perim}(P)=m$. Lemma 2.4 shows that
$r({P})=\frac{1}{2m}\cdot\underset{\bm{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\bm{e}_{j},\bm{u}\rangle|}.$ (5.1)
@@
$\min_{\begin{subarray}{c}\{\bm{e}_{1},\dotsc,\bm{e}_{m}\}\subset\mathbb{R}^{2}\end{subarray}}\underset{\bm{u}\in\mathbb{S}^{1}}{\max}$ $\,\sum_{j=1}^{m}{|\langle\bm{e}_{j},\bm{u}\rangle|}$ (5.2)
${\rm s.t.}$ $\,\,\sum_{j=1}^{m}\bm{e}_{j}=\bm{0},\,\,\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}=m.$
As shown in [3, Proposition 3], for any $\{\bm{e}_{1},\dotsc,\bm{e}_{m}\}\subset\mathbb{S}^{1}$, we have
$\underset{\bm{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\bm{e}_{j},\bm{u}\rangle|}=\max_{\bm{\varepsilon}\in\{\pm 1\}^{m}}\bigg\|\sum_{j=1}^{m}\varepsilon_{j}\cdot\bm{e}_{j}\bigg\|_{2}.$
Hence, if we assume that each $\|\bm{e}_{j}\|=1$, then (5.2) can be rewritten as
$\min_{\begin{subarray}{c}\{\bm{e}_{1},\dotsc,\bm{e}_{m}\}\subset\mathbb{S}^{1}\end{subarray}}\max_{\bm{\varepsilon}\in\{\pm 1\}^{m}}$ $\bigg\|\sum_{j=1}^{m}\varepsilon_{j}\cdot\bm{e}_{j}\bigg\|_{2}$ (5.3)
${\rm s.t.}$ $\,\,\sum_{j=1}^{m}\bm{e}_{j}=\bm{0}.$

### v2

@@ 3.1. The relationship between tight frames and convex polygons in $\mathbb{R}^{2}$
(ii) If $\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfies (1.11), by Theorem 1.3, there exists an optimal polygon $P\in{\mathcal{E}}_{m}$ with the edge vectors $\{\varepsilon_{j}(\cos(\frac{j}{m}\pi),\sin(\frac{j}{m}\pi))^{T}:j\in[m]\}$. Note that for $j\in[m]$ we have
$f({\boldsymbol{\nu}}_{j})=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}\quad\text{and}\quad f({\boldsymbol{\nu}}_{j}^{\perp})=-(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}.$ (4.13)
@@
Hence, the edge set of $P$ can be rewritten as $\{f({\boldsymbol{\nu}}_{j}):j\in I\}\cup\{f({\boldsymbol{\nu}}_{j}^{\perp}):j\in[m]\setminus I\}$ where $I=\{j\in[m]:\varepsilon_{j}=1\}$, and such a polygon $P$ is unique up to translation. Then it follows from Theorem 1.2 that the set $\{{\boldsymbol{\nu}}_{j}:j\in I\}\cup\{{\boldsymbol{\nu}}_{j}^{\perp}:j\in[m]\setminus I\}$ is an optimal tight frame.
@@
(iii) Let ${\mathcal{A}}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\in\mathcal{T}_{m}$. As established in the proof of (i), $\mathcal{A}$ is an optimal tight frame. Hence, by Theorem 1.2, there is an optimal strictly convex $m$-gon $P$ with the edge set $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$. According to Theorem 1.3, there exists a vector $\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfying (1.11), such that after a proper scaling and rotation, the edge set $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ equals $\{(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}:j\in I\}\cup\{-(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}:j\in[m]\setminus I\}$, where $I=\{j\in[m]:\varepsilon_{j}=1\}$. By (4.13), this edge set can be rewritten as $\{f({\boldsymbol{\nu}}_{j}):j\in I\}\cup\{f({\boldsymbol{\nu}}_{j}^{\perp}):j\in[m]\setminus I\}$. Note that the function $f(\cdot)$ is invertible on $\mathcal{T}$. It follows that ${\mathcal{A}}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}=\{{\boldsymbol{\nu}}_{j}:j\in I\}\cup\{{\boldsymbol{\nu}}_{j}^{\perp}:j\in[m]\setminus I\}$, up to scaling and rotation. This completes the proof. ∎
@@ 5. Discussion and Future Work
We conclude this paper by highlighting an intriguing connection between the perimeter-maximizing polygon problem and a fundamental question in discrepancy theory. Assume that $E=\{\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}\}$ is the edge set of a convex $m$-gon ${P}\subset\mathbb{R}^{2}$ with fixed perimeter $\mathrm{perim}(P)=m$. Lemma 2.4 shows that
$r({P})=\frac{1}{2m}\cdot\underset{\boldsymbol{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|}.$ (5.1)
@@
$\min_{\begin{subarray}{c}\{\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}\}\subset\mathbb{R}^{2}\end{subarray}}\underset{\boldsymbol{u}\in\mathbb{S}^{1}}{\max}$ $\,\sum_{j=1}^{m}{|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|}$ (5.2)
${\rm s.t.}$ $\,\,\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0},\,\,\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}=m.$
As shown in [3, Proposition 3], for any $\{\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}\}\subset\mathbb{S}^{1}$, we have
$\underset{\boldsymbol{u}\in\mathbb{S}^{1}}{\max}\,\sum_{j=1}^{m}{|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|}=\max_{\boldsymbol{\varepsilon}\in\{\pm 1\}^{m}}\bigg\|\sum_{j=1}^{m}\varepsilon_{j}\cdot\boldsymbol{e}_{j}\bigg\|_{2}.$
Hence, if we assume that each $\|\boldsymbol{e}_{j}\|=1$, then (5.2) can be rewritten as
$\min_{\begin{subarray}{c}\{\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}\}\subset\mathbb{S}^{1}\end{subarray}}\max_{\boldsymbol{\varepsilon}\in\{\pm 1\}^{m}}$ $\bigg\|\sum_{j=1}^{m}\varepsilon_{j}\cdot\boldsymbol{e}_{j}\bigg\|_{2}$ (5.3)
${\rm s.t.}$ $\,\,\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}.$